OpenStudy (anonymous):

Solve 3x - 4 ≤ 2 or 2x + 11 ≥ -1. {x | x ≤ 2} {x | -6 ≤ x ≤ 2} {all reals}

OpenStudy (akashdeepdeb):

Which one do you think would be the right answer after solving both the equations? :)

OpenStudy (anonymous):

3x - 4 <= 2 3x <= 2 + 4 3x <= 6 x <= 2 2x + 11 >= -1 2x >= -1-11 2x >= -12 x >= -6 -6 >= x <= 2